# Reliability-Latency-Rate Tradeoff in Low-Latency Communications with Finite-Blocklength Coding

### v3

@@ IV-B EC Approximation with Channel-Gain-Based Power Allocation
V Numerical Results
In this section, we conduct numerical calculations to validate the theoretical results and conclusions in this paper. Moreover, some characteristics of the statistical-delay performance of low-latency systems in the FBL regime will be explained more visually by the numerical results.
First, we validate the conclusions of high-SNR slopes of EC in the FBL regime in Figs. 5, 5, and 6. In these figures, we set $N=512$, $\epsilon=10^{-5}$, and $T=1$. In Fig. 5, $\mathbb{E}_{|h|^{2}}\{|h[n]|^{2}\}=1$. In Fig. 5, $f(x)$ satisfies Erlang distribution with $\kappa=2$ and $\mathbb{E}_{|h|^{2}}\{|h[n]|^{2}\}=2$. In Fig. 6, $\Omega=1$. For the Rayleigh fading case, $\frac{1}{T\log_{2}e}\approx 0.6931$. From Fig. 5, we find that the slopes of $\Lambda(\gamma)$ for $\theta N\leq 0.6931$, i.e. $\theta\leq 0.0014$, are nearly identical and close to 1, while the slopes of $\Lambda(\gamma)$ for $\theta>0.0014$ are smaller than 1. Besides, for those $\theta>0.0014$, a larger $\theta$ results in a smaller slope of $\Lambda(\gamma)$. The slope of $\Lambda(\gamma)$ for $\theta=0.004$ is $0.338$ in Fig. 5, while $\frac{1}{\theta NT\log_{2}e}$ with $\theta=0.004$ is 0.3385. The above results validate the accuracy of Corollary 1. Then we analyze the numerical results of the case in which $f(x)$ satisfies the Erlang distribution. In Fig. 5, $\frac{\kappa}{T\log_{2}e}\approx 1.3863$. We can find that with $\theta=0.002$, $\theta N=1.024$ is larger than $\frac{1}{T\log_{2}e}\approx 0.6931$ in Fig. 5 while it is smaller than $\frac{\kappa}{T\log_{2}e}\approx 1.3863$ in Fig. 5. Accordingly, the slope of $\Lambda(\gamma)$ for $\theta=0.002$ in Fig. 5 is different from the slope of ergodic capacity, while the slope of $\Lambda(\gamma)$ for $\theta=0.002$ is nearly identical to the slope of ergodic capacity in Fig. 5. This result reveals the difference between the performance of $s_{\infty}$ with and without the spatial or frequency diversity in Rayleigh fading channels, which also validates Corollary 1.
@@ V Numerical Results
In Fig. 6, we present the numerical results of $s_{\infty}$ in the Nakagami-$m$ fading channel. In Fig. 6(a), $\frac{m}{TN\log_{2}e}\approx 0.3466$, while in Fig. 6(b), $\frac{m}{TN\log_{2}e}\approx 1.3836$. The slopes of $\Lambda(\gamma)$ in Figs. 6(a) and 6(b) validates the accuracy of Theorem 5. Besides, by comparing the results in Figs. 5 and 6(b), we find that the slopes of $\Lambda(\gamma)$ is nearly identical with a same $\theta$. For example, when $\theta=0.004$, i.e., $\theta N=2.0480$, the slope of $\Lambda(\gamma)$ in Fig. 5 is $0.6731$, while the slope of $\Lambda(\gamma)$ in Fig. 6(b) is $0.6704$. This result also validates the accuracy of Theorem 5 and Corollary 1, since $s_{\infty}$ with $\kappa=2$ and $m=2$ are identical as shown in Theorem 5 and Corollary 1.
@@
Furthermore, we validate the accuracy of the approximation for EC by numerical calculations. First, we present the results of the EC approximation with the fixed-power allocation in Fig. 7. In this figure, $\gamma=10$ and $\epsilon=10^{-3}$. The channel model is set as the Rayleigh fading channel with $\mathbb{E}_{|h|^{2}}\{|h[n]|^{2}\}=1$. The numerical result in this figure is calculated based on the definition of the EC in Eq. (16), while the approximation result is obtained from Eq. (71). From Fig. 7, we find that the approximation results match well with the numerical results. The difference between the numerical and approximation results with $N=100$ and $T=1$ is slightly more evident when $\theta$ is small compared to the other three pairs. This is reasonable since the approximation error of Laplace’s method is $O\left(\frac{1}{\theta^{2}T^{2}N}\right)$. From these results, we see that the approximation in Lemma 2 satisfactorily approximates the EC in the low-latency scenarios.
Fig. 8 presents the result of EC approximation with the water-filling (WF) policy [46]. To test the performance of the EC approximation in different channel models, in Fig. 8 the channel gain satisfies $|h[n]|^{2}=|h_{1}[n]|^{2}+|h_{2}[n]|^{2}$, where $|h_{i}[n]|^{2}$ follows an exponential distribution with expectation 1. The approximation for the WF policy we adopted here is based on Eq. (77). From Fig. 8, we can see that the approximation and numerical results of the WF policy match well, which validates our analysis in Section III-C. Note that, for the water-filling policy with the average SNR $\mathbb{E}_{|h|^{2}}\{\gamma\}=10$, [^10] With some abuse of notations, we use $\gamma$ to denote the SNR of the water-filling policy, which should be a random variable related to $|h[n]|^{2}$. $\mathbb{E}_{{|h|^{2}}}\{s[n]\}\approx 3932$, while for the fixed-power allocation policy with $\gamma=10$, $\mathbb{E}_{|h|^{2}}\{s[n]\}\approx 3923$. This result indicates that the transmission optimization based on the infinite blocklength assumption can provide help for improving the average transmission rate in the FBL regime under some parameter settings. However, the result in Fig. 8 shows that, though the WF policy has a larger average transmission rate compared with the fixed-power allocation scheme, it does not improve the EC. Since the EC represents the maximum throughput under the QVP or DVP constraints, the poorer performance of the WF policy on the EC demonstrates that there is a difference between maximizing the average transmission rate and maximizing the throughput under statistical QoS constraints.
@@ VI Conclusion
In this paper, we mainly focused on characterizing the reliability-latency-rate tradeoff for low-latency systems with FBC. Based on the previous conclusions on the gain-conservation property for low-latency systems with deterministic arrival, we defined the reliability gain, real-time gain, and service-rate gain for the low-latency systems with random arrivals. EB and EC were adopted to establish the connection between the gain-conservation equation and the influence of random arrivals. By considering transmissions over an AWGN channel and a Nakagami-$m$ fading channel, we derived the service-rate gains in wired and wireless communication systems. Then, the revised gain-conservation equations were derived by taking service-rate, reliability, and real-time gains into account. Moreover, an EC approximation method was proposed to obtain the analytical expression of EC for some channel-gain-based scheduling policies in the FBL regime. The conclusions proposed in this paper characterize the fundamental tradeoff between the reliability, latency, and throughput for low-latency systems with FBC, which will provide instructions on the design of efficient scheduling schemes for the low-latency systems. Future promising research directions include obtaining more accurate approximations of QVP and DVP with the small threshold, relaxing the assumptions of the proposed analysis through tighter bounds or exact expression of the FBL coding rate, and designing efficient scheduling schemes based on the revised gain-conservation equations and proposed EC approximation method for low-latency communication systems.
@@ Appendix A Proof of Theorem 2
According to Eqs. (30) and (34), we have
${\color[rgb]{0,0,0}g_{u}(|h|^{2}\gamma,\epsilon)=\frac{c_{1}}{V\left(|h|^{2}\gamma\right)}+\frac{3}{2}\log_{2}V\left(|h|^{2}\gamma\right)-\log_{2}K\left(|h|^{2}\gamma\right),}\vspace{-1mm}$ (79)
and
@@
$\quad\quad\quad+\sqrt{NV\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{N}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon)\right],$
Since $|h|^{2}$ and $\gamma$ are independent, the product $|h|^{2}\gamma$ can be zero, finite, or approach positive infinity. We need to consider all possible cases of $|h|^{2}\gamma$. The essence of this proof is to demonstrate that $G(N,|h|^{2}\gamma,\epsilon)$ is bounded with the arbitrary value of $|h|^{2}\gamma$. We divide the proof into four parts. We let $\vartheta>0$ denote an arbitrarily small constant. The first part is the proof with $|h|^{2}\gamma\to 0$, i.e., $|h|^{2}\gamma\in[0,\vartheta)$. The second part is the proof with a bounded $\Psi(\gamma)$ and $|h|^{2}\gamma\in[\vartheta,+\infty)$. The third part is the proof with $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$ and $|h|^{2}\gamma\in[\vartheta,+\infty)$. Finally, we will summarize the proof in the fourth subsection. To use these bounds, we assume that
@@
The validity of Eq. (81) is provided in the fourth part.
@@
The analysis of $|h|^{2}\gamma\to 0$ is based on the physical significance. The proof in this subsection is to demonstrate that $G(N,|h|^{2}\gamma,\epsilon)$ is bounded when $|h|^{2}\gamma\to 0$. With $|h|^{2}\gamma\in\to 0$, it is obvious that $\tilde{R}^{\epsilon}_{N}$ is bounded according to Eq. (9). Since we can not obtain an unbounded rate with a finite received SNR, the coding rate $\tilde{R}^{\epsilon}_{N}+G(N,|h|^{2}\gamma,\epsilon)$ is also bounded as $|h|^{2}\gamma\to 0$. Thus, $G(N,|h|^{2}\gamma,\epsilon)$ must be bounded when $|h|^{2}\gamma\to 0$. This result implies $|G(N,|h|^{2}\gamma,\epsilon)|\leq u_{1}$ for $|h|^{2}\gamma\to 0$, where $u_{1}$ is a positive constant.
@@
We use the upper and lower bounds to prove that $G(N,|h|^{2}\gamma,\epsilon)$ is bounded as $|h|^{2}\gamma\to+\infty$. We start the analysis with the upper bound $g_{u}(|h|^{2}\gamma,\epsilon)$. According to Eq. (31), $K\left(|h|^{2}\gamma\right)$ is finite for a given finite $|h|^{2}\gamma\in[\vartheta,+\infty)$. As $|h|^{2}\gamma\to+\infty$, $\frac{\gamma}{1+\gamma}\to 1$ and
@@
Thus, $K\left(|h|^{2}\gamma\right)$ is also bounded when $|h|^{2}\gamma\in[\vartheta,+\infty)$.
Moreover, we have $\left(1-\frac{1}{(1+\vartheta)^{2}}\right)\left(\log_{2}e\right)^{2}\leq V\left(|h|^{2}\gamma\right)\leq\left(\log_{2}e\right)^{2}$ when $|h|^{2}\gamma\in[\vartheta,+\infty)$. Thus, we have obtained that $K\left(|h|^{2}\gamma\right)$ and $V\left(|h|^{2}\gamma\right)$ is bounded for $|h|^{2}\gamma\in[\vartheta,+\infty)$.
Then, we will analyze $b_{1}$. First, we want to show that $b_{1}<1-\epsilon$, i.e.,
@@
${\color[rgb]{0,0,0}\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{(2K\left(|h|^{2}\gamma\right))^{2}+V^{3}\left(|h|^{2}\gamma\right)(1-\epsilon)^{2}}}<1.}$ (84)
Eq. (84) is true since $\left(1-\frac{1}{(1+\vartheta)^{2}}\right)\left(\log_{2}e\right)^{2}\leq V\left(|h|^{2}\gamma\right)$ with $|h|^{2}\gamma\in[\vartheta,+\infty)$. Thus, $01-\epsilon-b_{1}>$ 0, which indicates $\min_{y\in\left[1-\epsilon-b_{1},1-\epsilon\right]}\frac{{\rm d}Q^{-1}(y)}{{\rm d}y}$ is finite. Therefore, $c_{1}$ is bounded for $|h|^{2}\gamma\in[\vartheta,+\infty)$ according to Eq. (32).
We have proved that for $|h|^{2}\gamma\in[\vartheta,+\infty)$, every term of $g_{u}\left(|h|^{2}\gamma,\epsilon\right)$ shown in Eq. (79) is finite. Thus, $|g_{u}\left(|h|^{2}\gamma,\epsilon\right)|\leq u_{2}$ for $|h|^{2}\gamma\in[\vartheta,+\infty)$, where $u_{2}$ is a positive constant.
Then, we will analyze $g_{l}(N,|h|^{2}\gamma,\epsilon)$. We consider two cases. Similar to the analysis of $g_{u}\left(|h|^{2}\gamma,\epsilon\right)$, every term of $g_{l}(N,|h|^{2}\gamma,\epsilon)$ is finite for $|h|^{2}\gamma\in[\vartheta,+\infty)$. This result implies $\left|g_{l}(N,|h|^{2}\gamma,\epsilon)\right|\leq u_{3}$, where $u_{3}$ is a positive constant.
In summary, we have proved that $|g_{u}\left(|h|^{2}\gamma,\epsilon\right)|$ and $|g_{l}(N,|h|^{2}\gamma,\epsilon)|$ are all bounded for $|h|^{2}\gamma\in[\vartheta,+\infty)$ and a bounded $\Psi(\gamma)$. Since $g_{l}(N,|h|^{2}\gamma,\epsilon)\leq G(N,|h|^{2}\gamma,\epsilon)\leq g_{u}(|h|^{2}\gamma,\epsilon)$, we have $|G(N,|h|^{2}\gamma,\epsilon)|\leq\max\{u_{2},u_{3}\}$ for $|h|^{2}\gamma\in[\vartheta,+\infty)$.
Combining the results from the last subsection with this subsection, we have proved that
holds with a given $\epsilon$, a bounded $\Psi(\gamma)$, and a given $N$ satisfying Eq. (81). With a bounded $\Psi(\gamma)$, Eq. (85) ensures Eq. (22) is satisfied. Thus, to satisfy the condition in Eq. (22) with a given $\epsilon$ and a bounded $\Psi(\gamma)$, we only need to ensure that Eq. (81) is met.
Since the upper bound is irrelevant to $N$, the only difference between this part and the last part lies in the analysis of the lower bound $g_{l}(N,|h|^{2}\gamma,\epsilon)$. For $|h|^{2}\gamma\in[0,\vartheta)$, the proof is the same as the proof in Subsection A. Thus, we only need to focus on $|h|^{2}\gamma\in[\vartheta,+\infty)$. With $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, there are two unbounded terms in $g_{l}(N,|h|^{2}\gamma,\epsilon)$. The unbounded terms are $-\frac{1}{2}\log_{2}\Psi(\gamma)$ and $\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon(\gamma)+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon(\gamma))\right]$. Since $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, we have
Then, we turn to the term $\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon)\right]$. By performing a Taylor-series expansion of the inverse Gaussian Q-function around $1-\epsilon$, we obtain
@@
Therefore, for the lower bound $g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))$, any $\Psi(\gamma)$ that satisfies $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$ can ensure $\lim\limits_{\gamma\to+\infty}\frac{g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}=0$. Since the upper bound is bounded, we have $\lim\limits_{\gamma\to+\infty}\frac{g_{u}(|h|^{2}\gamma,\epsilon)}{\Psi(\gamma)}=0$. Since we use the upper and lower bounds, we need to ensure that Eq. (81) holds. Thus, we have proved that, with a $\Psi(\gamma)$ satisfying Eq. (81) and $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, Eq. (22) holds.

### v4

@@ V Numerical Results
In Fig. 6, we present the numerical results for $s_{\infty}$ in the Nakagami-$m$ fading channel. In Fig. 6(a), $\frac{m}{TN\log_{2}e}\approx 0.3466$, while in Fig. 6(b), $\frac{m}{TN\log_{2}e}\approx 1.3836$. The slopes of $\Lambda(\gamma)$ in Figs. 6(a) and 6(b) validate the accuracy of Theorem 5. Besides, by comparing the results in Figs. 5 and 6(b), we find that the slopes of $\Lambda(\gamma)$ are nearly identical for the same value of $\theta$. For example, when $\theta=0.004$, i.e., $\theta N=2.0480$, the slope of $\Lambda(\gamma)$ in Fig. 5 is $0.6731$, while it is $0.6704$ in Fig. 6(b). This finding further validates the accuracy of Theorem 5 and Corollary 1, as $s_{\infty}$ with $\kappa=2$ and $m=2$ are shown to be identical in both Theorem 5 and Corollary 1.
@@
Additionally, we validate the accuracy of the EC approximation through numerical calculations. In Fig. 7, we present the results of the EC approximation for the fixed-power allocation, where $\gamma=10$ and $\epsilon=10^{-3}$. The channel is modeled as a Rayleigh fading channel with $\mathbb{E}_{|h|^{2}}\{|h[n]|^{2}\}=1$. The results with the legend “numerical” are calculated based on the definition of the EC in Eq. (16), while the results with the legend “approximation” are obtained from Eq. (72). As shown in Fig. 7, we find that the approximation results match well with the numerical results. However, the difference between the numerical and approximation results with $N=100$ and $T=1$ is slightly more evident when $\theta$ is small compared to the other three pairs. This is reasonable since the approximation error of Laplace’s method is $O\left(\frac{1}{\theta^{2}T^{2}N}\right)$. From these results, we see that the approximation in Lemma 2 satisfactorily approximates the EC in the low-latency scenarios.
Fig. 8 illustrates the results of EC approximation for the water-filling (WF) policy [46]. To evaluate the performance of the EC approximation in different channel models, the channel gain in Fig. 8 is set as $|h[n]|^{2}=|h_{1}[n]|^{2}+|h_{2}[n]|^{2}$, where $|h_{i}[n]|^{2}$ follows an exponential distribution with expectation 1. The approximation for the WF policy used here is based on Eq. (78). As seen in Fig. 8, the approximation results closely matches the numerical results, which validates our analysis in Section III-C. Note that, for the WF policy with the average SNR $\mathbb{E}_{|h|^{2}}\{\gamma\}=10$, [^10] For simplicity, we use $\gamma$ to denote the SNR of the WF policy, which should be a random variable related to $|h[n]|^{2}$. the average transmission rate $\mathbb{E}_{{|h|^{2}}}\{s[n]\}\approx 3932$, while for the fixed-power allocation policy with $\gamma=10$, $\mathbb{E}_{|h|^{2}}\{s[n]\}\approx 3923$. This result indicates that the transmission optimization under the infinite blocklength assumption can improve the average transmission rate in the FBL regime under certain parameter settings. However, the result in Fig. 8 also shows that, although the WF policy leads to a larger average transmission rate compared to the fixed-power allocation scheme, it does not improve the EC. Since the EC represents the maximum throughput under the QVP or DVP constraints, the relatively poorer performance of the WF policy on the EC highlights the distinction between maximizing the average transmission rate and maximizing throughput under statistical QoS constraints.
@@ VI Conclusion
In this paper, we mainly focused on characterizing the fundamental tradeoff between reliability, latency, and rate for low-latency systems operating under FBC. Based on the previous conclusions on the gain-conservation properties for low-latency systems with deterministic arrival, we extended the definitions of the reliability gain, real-time gain, and service-rate gain for the low-latency systems with random arrivals. By employing EB and EC, we established the connection between the gain-conservation equation and the influence of random arrivals. Our analysis, conducted over both AWGN and Nakagami-$m$ fading channels, provided insights into service-rate gains for both wired and wireless systems. Then, the revised gain-conservation equations were derived by incorporating service-rate, reliability, and real-time gains. Additionally, an EC approximation method was proposed to derive an analytical expression of EC in systems with channel-gain-based scheduling policies in the FBL regime. The conclusions proposed in this paper offer a comprehensive characterization of the fundamental tradeoff between the reliability, latency, and throughput for low-latency systems with FBC, which will provide instructions for designing efficient scheduling schemes for the low-latency systems. Future research directions include refining the approximations of QVP and DVP with small thresholds, relaxing the assumptions of the proposed analysis through tighter bounds or exact expression of the FBC rate, and developing advanced scheduling schemes based on the revised gain-conservation equations and the proposed EC approximation method for low-latency communication systems.
@@ Appendix A Proof of Theorem 2
According to Eqs. (31) and (35), we derive Eqs. (80) and (81), which are provided at the bottom of the next page.
$g_{u}(|h|^{2}\gamma,\epsilon)=\frac{c_{1}}{V\left(|h|^{2}\gamma\right)}+\frac{3}{2}\log_{2}V\left(|h|^{2}\gamma\right)-\log_{2}K\left(|h|^{2}\gamma\right).$ (80)
@@
$\quad\quad\quad+\sqrt{NV\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{N}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon)\right].$
Since $|h|^{2}$ and $\gamma$ are independent, the product $|h|^{2}\gamma$ can be zero, finite, or approach positive infinity. Hence, we have to consider all possible cases of $|h|^{2}\gamma$. The essence of this proof is to demonstrate that $G(N,|h|^{2}\gamma,\epsilon)$ remains bounded regardless of the value of $|h|^{2}\gamma$. The proof is divided into four parts. We let $\vartheta>0$ denote an arbitrarily small constant. The first part is the proof for $|h|^{2}\gamma\to 0$, i.e., $|h|^{2}\gamma\in[0,\vartheta)$. The second part is the proof for a bounded $\Psi(\gamma)$ and $|h|^{2}\gamma\in[\vartheta,+\infty)$. The third part is the proof for $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$ and $|h|^{2}\gamma\in[\vartheta,+\infty)$. Finally, we will summarize the proof in the fourth subsection. To apply the bounds shown in Eqs. (80) and (81), we assume that
@@
The validity of Eq. (82) is established in the fourth part.
@@
The analysis of the case where $|h|^{2}\gamma\to 0$ is grounded in its physical significance. The proof in this subsection is to demonstrate that $G(N,|h|^{2}\gamma,\epsilon)$ remains bounded as $|h|^{2}\gamma\to 0$. When $|h|^{2}\gamma\in\to 0$, it is clear that $\tilde{R}^{\epsilon}_{N}$ is bounded according to Eq. (9). Since we can not obtain an unbounded rate with a finite received SNR, the coding rate $\tilde{R}^{\epsilon}_{N}+G(N,|h|^{2}\gamma,\epsilon)$ must also be bounded as $|h|^{2}\gamma\to 0$. Consequently, $G(N,|h|^{2}\gamma,\epsilon)$ must remain bounded when $|h|^{2}\gamma\to 0$. This result implies that $|G(N,|h|^{2}\gamma,\epsilon)|\leq u_{1}$ when $|h|^{2}\gamma\to 0$, where $u_{1}$ is a positive constant.
@@
We use the upper and lower bounds to demonstrate that $G(N,|h|^{2}\gamma,\epsilon)$ remains bounded as $|h|^{2}\gamma\to+\infty$. We start the analysis with the upper bound $g_{u}(|h|^{2}\gamma,\epsilon)$. According to Eq. (32), $K\left(|h|^{2}\gamma\right)$ is finite for a given finite $|h|^{2}\gamma\in[\vartheta,+\infty)$. As $|h|^{2}\gamma\to+\infty$, $\frac{\gamma}{1+\gamma}\to 1$ and
@@
Thus, $K\left(|h|^{2}\gamma\right)$ remains bounded when $|h|^{2}\gamma\in[\vartheta,+\infty)$.
Moreover, we have
$\left(1-\frac{1}{(1+\vartheta)^{2}}\right)\left(\log_{2}e\right)^{2}\leq V\left(|h|^{2}\gamma\right)\leq\left(\log_{2}e\right)^{2},$
for $|h|^{2}\gamma\in[\vartheta,+\infty)$. This implies that both $K\left(|h|^{2}\gamma\right)$ and $V\left(|h|^{2}\gamma\right)$ are bounded for $|h|^{2}\gamma\in[\vartheta,+\infty)$.
Next, we analyze $b_{1}$. First, we want to show that $b_{1}<1-\epsilon$, i.e.,
@@
$\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{(2K\left(|h|^{2}\gamma\right))^{2}+V^{3}\left(|h|^{2}\gamma\right)(1-\epsilon)^{2}}}<1.$ (85)
Eq. (85) holds since $\left(1-\frac{1}{(1+\vartheta)^{2}}\right)\left(\log_{2}e\right)^{2}\leq V\left(|h|^{2}\gamma\right)$ with $|h|^{2}\gamma\in[\vartheta,+\infty)$. Hence, $01-\epsilon-b_{1}>$ 0 indicating that $\min_{y\in\left[1-\epsilon-b_{1},1-\epsilon\right]}\frac{{\rm d}Q^{-1}(y)}{{\rm d}y}$ is finite. Therefore, $c_{1}$ is bounded for $|h|^{2}\gamma\in[\vartheta,+\infty)$ according to Eq. (33).
Thus, we have proved that for $|h|^{2}\gamma\in[\vartheta,+\infty)$, every term of $g_{u}\left(|h|^{2}\gamma,\epsilon\right)$ shown in Eq. (80) is finite. Consequently, $|g_{u}\left(|h|^{2}\gamma,\epsilon\right)|\leq u_{2}$ for $|h|^{2}\gamma\in[\vartheta,+\infty)$, where $u_{2}$ is a positive constant.
@@
Then, we will analyze $g_{l}(N,|h|^{2}\gamma,\epsilon)$. We consider two cases. Following a similar approach to the analysis of $g_{u}\left(|h|^{2}\gamma,\epsilon\right)$, every term in $g_{l}(N,|h|^{2}\gamma,\epsilon)$ is finite for $|h|^{2}\gamma\in[\vartheta,+\infty)$. This result implies that $\left|g_{l}(N,|h|^{2}\gamma,\epsilon)\right|\leq u_{3}$, where $u_{3}$ is a positive constant.
In summary, we have proved that both $|g_{u}\left(|h|^{2}\gamma,\epsilon\right)|$ and $|g_{l}(N,|h|^{2}\gamma,\epsilon)|$ are bounded for $|h|^{2}\gamma\in[\vartheta,+\infty)$ and a bounded $\Psi(\gamma)$. Since $g_{l}(N,|h|^{2}\gamma,\epsilon)\leq G(N,|h|^{2}\gamma,\epsilon)\leq g_{u}(|h|^{2}\gamma,\epsilon)$, it follows that $|G(N,|h|^{2}\gamma,\epsilon)|\leq\max\{u_{2},u_{3}\}$ for $|h|^{2}\gamma\in[\vartheta,+\infty)$.
Combining the results from the last subsection with those of this subsection, we have proved that
holds for a given $\epsilon$, a bounded $\Psi(\gamma)$, and an $N$ satisfying Eq. (82). With a bounded $\Psi(\gamma)$, Eq. (86) ensures that Eq. (22) is satisfied. Therefore, to satisfy the condition in Eq. (22) with a given $\epsilon$ and a bounded $\Psi(\gamma)$, it is sufficient to ensure that Eq. (82) is fulfilled.
Since the upper bound is independent of $N$, the only difference between this subsection and the last subsection lies in the analysis of the lower bound $g_{l}(N,|h|^{2}\gamma,\epsilon)$. For $|h|^{2}\gamma\in[0,\vartheta)$, the proof is the same as that presented in Subsection A. Therefore, the focus is placed on $|h|^{2}\gamma\in[\vartheta,+\infty)$. When $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, two terms in $g_{l}(N,|h|^{2}\gamma,\epsilon)$ are unbounded. The unbounded terms are $-\frac{1}{2}\log_{2}\Psi(\gamma)$ and
$\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}$ $\Bigg{[}Q^{-1}(\epsilon(\gamma))+$
$Q^{-1}$ $\left(1-\epsilon(\gamma)+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)\Bigg{]}.$
Since $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, it follows that
Next, we consider the term shown in Eq. (A-C). By performing a Taylor-series expansion of the inverse Gaussian Q-function around $1-\epsilon$, we obtain Eq. (88), which appears at the bottom of this page. Consequently, for the lower bound $g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))$, any $\Psi(\gamma)$ that satisfies $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$ can ensure $\lim\limits_{\gamma\to+\infty}\frac{g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}=0$. Given that the upper bound is also bounded, we have $\lim\limits_{\gamma\to+\infty}\frac{g_{u}(|h|^{2}\gamma,\epsilon)}{\Psi(\gamma)}=0$. Since we use the upper and lower bounds, it is essential to verify that Eq. (82) holds. Thus, we have demonstrated that, with a $\Psi(\gamma)$ satisfying Eq. (82) and $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, Eq. (22) holds.